# SHaRe-SSM: An Oscillatory Spiking Neural Network for Target Variable Modeling in Long Sequences

### v1

@@ 4 Emperical Results
4.1 Long-range sequence classification
Table 1: Classification performance (mean and standard deviation across five training runs) for UEA time-series datasets using hyperparameter protocol from walker2024log . For convenience, some dataset names are abbreviated: Worms (EigenWorms), SCP1 (SelfRegulationSCP1), SCP2 (SelfRegulationSCP2), Ethanol (EthanolConcentration), Heartbeat, and Motor (MotorImagery). The top two models are bold and underlined, respectively. SNN implies a spiking neural network(Y/N).
Seq. Length/ Class 17,984/5 896/2 1,152/2 1,751/4 405/2 3,000/2
SNN Worms SCP1 SCP2 Ethanol Heartbeat Motor Avg
NRDE N $83.9\pm 7.3$ $80.9\pm 2.5$ $53.7\pm 6.9$ $25.3\pm 1.8$ $72.9\pm 4.8$ $47.0\pm 5.7$ $60.6$
NCDE N $75.0\pm 3.9$ $79.8\pm 5.6$ $53.0\pm 2.8$ $29.9\pm 6.5$ $73.9\pm 2.6$ $49.5\pm 2.8$ $60.2$
Log-NCDE N $85.6\pm 5.1$ $83.1\pm 2.8$ $53.7\pm 4.1$ $\bf 34.4\pm 6.4$ $75.2\pm 4.6$ $53.7\pm 5.3$ $64.3$
LRU N $87.8\pm 2.8$}} $82.6\pm 3.4$ $51.2\pm 3.6$ $21.5\pm 2.1$ $\bf 78.4\pm 6.7$ $48.4\pm 5.0$ $61.7$
S5 N $81.1\pm 3.7$ $\bf{89.9\pm 4.6}$} $50.5\pm 2.6$ $24.1\pm 4.3$ $77.7\pm 5.5$ $47.7\pm 5.5$ $61.8$
S6 N $85.0\pm 16.1$ $82.8\pm 2.7$ $49.9\pm 9.4$ $26.4\pm 6.4$ $76.5\pm 8.3$ $51.3\pm 4.7$ $62.0$
Mamba N $70.9\pm 15.8$ $80.7\pm 1.4$ $48.2\pm 3.9$ $27.9\pm 4.5$ $76.2\pm 3.8$ $47.7\pm 4.5$ $58.6$
LinOSS (IM) N $\bf{95.0\pm 4.4}$} $87.8\pm 2.6$}} $58.2\pm 6.9$}} $29.9\pm 0.6$}} $75.8\pm 3.7$ $60.0\pm 7.5$}} 67.8{67.8}
LinOSS (IMEX) N $80.0\pm 2.7$ $87.5\pm 4.0$}} $\bf{58.9\pm 8.1}$} $29.9\pm 1.0$ $75.5\pm 4.3$}} $57.9\pm 5.3$}} 65.0{{65.0}}
D-LinOSS N $93.9\pm 3.2$ $88.9\pm 3.0$ $58.6\pm 2.3$ $29.9\pm 0.6$ $75.8\pm 4.9$ $\bf 61.1\pm 2.0$ $\bf 68.0$
Ours (IM) Y $85.6\pm 8.1$ $82.6\pm 4.5$} $\bf 55.4\pm 6.9$} $30.1\pm 1.47$ $\bf 74.1\pm 5.9$} $\bf 59.6\pm 5.4$ $\bf 64.4$
Ours (IMEX) Y $\bf 85.6\pm 2.7$ $\bf 82.6\pm 4.2$ $53.0\pm 4.8$ $\bf 32.4\pm 4.6$ $73.2\pm 5.73$ $58.6\pm 7.3$ $64.2$
For classification experiments, we show SHaRe-SSM’s efficacy on the long-range sequential benchmark introduced in walker2024log ; rusch2025oscillatory with pre-selected random seeds for splitting the dataset (70/15/15 splits) and a fixed model hyperparameter range. It consists of the six longest sequence datasets pertaining to real-world applications from the University of East Anglia Multivariate Time Series Classification Archive (UEA-MTSCA) bagnall2018uea . To the best of our knowledge, SHaRe-SSM is the first second-order spiking SSM proposed. In addition, the recently proposed ANN-based second-order SSM (LinOSS) rusch2025oscillatory very clearly establishes its superiority over first-order SSMs. Moreover, presently proposed spiking first-order SSM’s performance pspikessm ; spikingssm ; stan2024learning is bounded by S4D’s performance. Hence, for these extremely long-range classification tasks, we compare our model performance only to the ANN-based State Space Modelss5 ; mamba ; rusch2025oscillatory (First and Second order) and Neural controlled ODEsnrde ; ncde ; walker2024log only. ANN-based LinOSS rusch2025oscillatory leverages second-order ODEs for long-range sequence learning and outperforms other SOTA models on such very long-range benchmark datasets. Similarly, SHaRe-SSM, on average, performs better than Mamba & S6 by $5.7\%$ & $2.3\%$ (See Table 1 respectively, while being only 3.6% lower than SOTA D-LinOSS. Our simplistic but fully spike-based model design (SHaRe-SSM), which leverages HRF neuronal dynamics, performs just $9.4\%$ lower than LinOSS for IM discretization on the 18k Eigenworms dataset (noteworthily (pspikessm, ) also observed a 9.4% performance drop on the Pathfinder dataset when compared to S4). Whereas for IMEX discretization, SHaRe-SSM outperforms LinOSS by $5.6\%$, while being extremely energy efficient (see Table 1). Moreover, we notice a trend that even though accuracy is the same, IMEX standard deviation is lower by $5.41\%$ and $0.32\%$ than IM for EigenWorms and SCP1, respectively. This strengthens the claim that IM introduces dissipative terms that can impact model performance. Similarly, we observe that for the Ethanol data set, the IMEX discretization on SHaRe-SSM outperforms all variants of LinOSS by $2.5\%$. While Log-NCDE outperforms our model, it comes with its own ODE computational overheads compared to LinOSS. On the 3k MotorImagery dataset, SHaRe-SSM achieves a performance of 59.6% (vs. 61.1% for D-LinOSS). Overall, our second-order spiking SSM outperforms first-order SSMs based on Mamba. Notably, a drop in performance with the gain in efficiency is ubiquitous across spiking SSMs for LRA tasks stan2024learning ; pspikessm ; spikingssm . To ensure a fair comparison, we followed the hyperparameter protocol from walker2024log ; However, constrained by computational resources and time, instead of grid search, we ran Bayesian search using Optuna akiba2019optuna for 15 runs and performed a narrow search.
@@ 4.1.1 Energy computation
Figure 3: Variation of energy computation of a LinOSS Block rusch2025oscillatory to the SHaRe-SSM Block for EigenWorms with respect to State (P), Hidden size (H).
In this section, we assess and compare the energy computation for LinOSS and SHaRe-SSM for similar hidden-size(H), state-size(P), and num-blocks(N) on the EigenWorms dataset. We estimate the theoretical energy consumption of our model based on prior works spikingssm ; pspikessm ; wu2025spikf . Accordingly, we assume that MAC and AC operations are implemented on a 45nm hardware horowitz20141 , where $E_{\text{MAC}}=4.6\text{pJ}$ and $E_{\text{AC}}=0.9\text{pJ}$. Notably, as discussed in  (spikingssm, ), the computational cost of multiplication of a floating-point weight by a binary activation number is represented as an addition-only operation. For ANNs, the theoretical energy consumption of a block $n$ is given by $4.6\text{pJ}\times\text{FLOPs}(n)$. For SNNs, the energy consumption for $n$ is given by $0.9\text{pJ}\times\text{SOPs}(n)$. Calculating theoretical energy consumption requires first calculating the synaptic operations, $\text{SOPs}(n)=f_{r}\times\text{FLOPs}(n)$, $f_{r}$ is the firing rate of the input spike train of the block/layer, $\text{FLOPs}(n)$ refers to the number of floating-point operations in layer $n$, equivalent to the number of multiply-and-accumulate (MAC) operations. SOPs denote the number of spike-based accumulate (AC) operations.
Since $\Omega$ is a diagonal matrix, it can be efficiently implemented using parallel scans with computations of order $\mathcal{O}(P\,\log(L))$, which is negligible. Also, D is $\mathcal{O}(HL)$, which is much smaller than $\mathcal{O}(LPH)$, i.e., computation for B & C matrices or even $\mathcal{O}(LH^{2})$, which is for the GLU layer. rusch2025oscillatory uses GeLU and GLU non-linearities, which we replace with a linear layer. And, yu2023metaformer described that GeLU consumes 14 FLOPs per operation. Moreover, GLU has twice as many FLOPs as a linear layer.
@@
From Figure 3, we observe that for different state-to-hidden ratios, we get different energy numbers. For the EigenWorms dataset, the average spike rate per time step per block is $0.32$. For a lower P/H ratio, we yield high energy efficiencies ($\sim 120\times$). Even for a high ratio, our model block is $\sim 20\times$ more energy-efficient than a LinOSS Block for EigenWorms. Our best performing model (as per Table 1) is $\textbf{73.2}\times$ more energy efficient than an equivalent LinOSS model.
4.2 Regression with spikes
None of the previously proposed spiking SSMs report performance on regression tasks stan2024learning ; pspikessm ; spikingssm . In fact, most of the spiking deep learning research is focused on classification zhou2023spikformer ; higuchi2024balanced . Recently, a transformer-based method wu2025spikf has been proposed for spike-based regression of time series. But this approach is significantly limited (and deviates from present hardware-supported featuresShresthaLoihi2 ) due to its quadratic sequence length dependence bottleneck and computationally intensive FFT computations pspikessm . Unlike a transformer, SSMs have emerged as a powerful solution to the problem of such quadratic dependence on sequence length. To the best of our knowledge, for the first time, we report the regression performance of a spiking SSM for extremely long sequences (up to 50k length). Apparently, a spike-based output limits the output range for regression tasks, making it a difficult task for spiking neurons. Previously, a spike-free LI layer was shown to work for some datasets Henkes2024-zk . However, it is not scalable to the size of very long sequences due to the oversmoothing effect of the moving average. Another approach wu2025spikf translates spikes back into continuous values using a Linear layer. However, this method also leads to limited expressivity about the time steps. To mitigate these issues, we propose a kernel-based regression with a learnable kernel convolved over time, initialized with a decaying LI filter of fixed size. To find the best parameters, we apply Bayesian hyperparameter akiba2019optuna search instead of the grid search with the same range of hyperparameters for a fair comparison. Subsequently, we ran a grid search on he best kernel size to ensure our model doesn’t lose its representation power, nor oversmooths the outputs(see Table 5).
4.2.1 PPG-DaLiA Dataset Evaluation
To model very long-range interactions for regression tasks, we consider a 49920-length sequence following the processing and hyperparameter protocol as proposed in walker2024log . PPG-DalLiA is a multivariate time series regression dataset to predict heart rate using data collected from a wrist-worn device from 15 individuals for 150 minutes, sampled at 128Hz. We report the best MSE in the order of $10^{-2}$ after applying Bayesian search for 15 runs, and applied the same hyperparameters for different discretizations. From Table 2, we observe that SHaRe-SSM models consistently outperform all first-order SSM variants, even after spiking in nature. Our best-performing model is SHaRe-SSM-IMEX, which outperforms Mamba by $0.0057$ MSE, proving the energy-conserving nature of IMEX-discretization and our second-order resonating neuron representation power. Moreover, despite being energy efficient, we are only $0.0392$ MSE lower than second-order SSMs.
Table 2: Mean and Standard Deviation reported for PPG-DaLiA dataset using hyperparameter protocol from walker2024log for the mean-squared error (MSE) across five training runs. For a clear differentiation, spiking neural networks are kept separately.
Model SNN MSE $\times 10^{-2}$
NRDE N $9.90\pm 0.97$
NCDE N $13.54\pm 0.69$
Log-NCDE N $9.56\pm 0.59$}}
LRU N $12.17\pm 0.49$
S5 N $12.63\pm 1.25$
S6 N $12.88\pm 2.05$
Mamba N $10.65\pm 2.20$
LinOSS-IMEX N $7.5\pm 0.46$}}
LinOSS-IM N $6.4\pm 0.23$}}
D-LinOSS N $\bf{6.16\pm 0.73}$
Ours (IM) Y $11.27\pm 0.67$
Ours (IMEX) Y $\bf{10.08\pm 0.46}$
4.2.2 Weather Dataset Evaluation
To test our model’s performance, we follow the hyperparameter protocol of rusch2025oscillatory on a time series dataset that utilizes domain-specific preprocessing. We consider the most challenging task of 720-timestep prediction (hours) for the weather dataset from informer . We report the best Mean Absolute Error with mean and standard deviation for our model on IM and IMEX discretizations. Also, we perform Bayesian search instead of grid search with the same hyperparameter protocol and report performance for both discretizations. SHaRe-SSM outperforms ANN transformer-based modelsinformer ; logtrans ; reformer (Informer by $0.145$). RNN-based modelslstma ; lstnet such as LSTnet by $0.171$ and competitive performance with S4, with just a $0.008$ drop. From Table 3, we notice that IMEX has a deviation lower than $0.003$ compared to IM because of its energy-conserving properties.
Table 3: Mean and standard deviation (across five runs) of mean-absolute error (MAE) on the weather dataset while predicting future 720 timesteps. Hyperparameter protocol from rusch2025oscillatory
Model SNN MAE
Informer N $0.731$
Informer N $0.741$
LogTrans N $0.773$
Reformer N $1.575$
LSTMa N $1.109$
LSTnet N $0.757$
S4 N $0.578$
LinOSS-IMEX N $0.508$
LinOSS-IM N $0.528$
D-LinOSS N $0.486$
Ours (IMEX) Y $\bf{0.586\pm 0.001}$
Ours (IM) Y $0.586\pm 0.004$
4.3 Ablation Studies
4.3.1 Heterogeneity Analysis
Given the established sensitivity of SSMs to initializations, SHaRe-SSM is incorporated with heterogeneities for components - $\Omega$, B, C, D matrices and $\Delta t$, of the HRF-SSM, and the thresholds for the C ($\theta_{C}$), D ($\theta_{D}$) matrices and, the encoder ($\theta_{\text{encoder}}$). We did a detailed ablation study of the effect of each component’s heterogeneity on performance by making them homogeneous (initialized by a constant value across all neurons).
Table 4: Ablation study - Studying the effect of making each component fixed (homogeneous) on the EigenWorms dataset using IMEX with the best parameters for five seeds.(Here, $x^{*}=x^{-1/2}$)
Heterogeneity Value Accuracy
$\theta_{\text{encoder}}$ $\mathcal{U}(0,1]$ 1.0 $85.0\pm 7.2$
$\Omega$ $\mathcal{U}(0,1]$ 1.0 $82.2\pm 3.3$
B $\mathcal{U}(-{H}^{*},{H}^{*})$ 0.0 $85.0\pm 4.5$
C $\mathcal{U}(-{P}^{*},{P}^{*})$ 0.0 $85.0\pm 4.5$
D $\mathcal{N}(0,1)$ 0.0 $80.5\pm 4.6$
$\Delta t$ $\mathcal{U}(0,1]$ 1.0 $82.8\pm 2.7$
$\theta_{C}$ $\mathcal{U}(0,1]$ 1.0 $83.3\pm 3.9$
$\theta_{D}$ $\mathcal{U}(0,1]$ 1.0 $85.0\pm 4.8$
$\Omega$,B,C,D As Above As Above $85.5\pm 4.1$
$\theta_{C}$, $\theta_{D}$ As Above As Above $83.9\pm 5.1$
All As Above As Above $85.5\pm 4.1$
Only SSM Default Default $87.2\pm 3.3$
From Table 4, we observe that homogeneous initialization for neuronal thresholds leads to decreased performance. For homogeneous big step size, we see performance drop from 85.6 by 2.8% (See Table 4). Further, it drops down by 3.4% and 4.5% if the $\Omega$ matrix and D matrix are initialized with frequencies to be homogeneously kept as one and zero, respectively, showing the importance of heterogeneous states in these networks. It drops by 5% if D is initialized with zero. This implies, it is desirable to mix the HRF output with the input spikes. Homogeneity in thresholds in the Encoder layer doesn’t drop the performance much (only 0.6%), but for the thresholds of C and D, it drops by 1.7%, implying the importance of heterogeneous initialization for HRF neurons. We also noticed that complete homogenization of the model didn’t alter the mean accuracy much; rather, the deviation nearly doubled. Interestingly, if we just consider the SSM layer in a SHaRe-SSM block, we observe a boost in model performance for EigenWorms by $1.6\%$, possibly because of higher representation power without the linear layer.
4.3.2 Kernel Size for Regression
For very long-range sequences, to avoid oversmoothing, but at the same time increasing the representation power of a linear decoder, we introduce convolving a filter. This can be thought of as similar to an LI filter with learnable parameters. We investigate kernel size on the best performing model from Table 3. From Table 5, we observe that for the 50k PPG-DaLiA dataset, a kernel size of 64 outperforms the 32 and 128 filters by an error of 0.005.
Table 5: Ablation on kernel size for regression task
Model Kernel Size MSE $\times 10^{-2}$
IMEX 32 $10.62\pm 0.59$
64 $\bf{10.08\pm 0.46}$
128 $10.69\pm 0.45$
@@ 5 Discussion
In this work, we investigate the plausibility of an integration of second-order spiking neurons in a state space model (SSM) framework. As part of it, we propose an energy-efficient second-order spiking SSM: SHaRe-SSM using harmonic resonate and fire neurons. In addition, we design a learnable encoder, decoder, and a compatible parallel scan method. SHaRe-SSM is fully spike-based without any GeLU or GLU. And, SHaRe-SSM is uniquely designed to achieve energy-efficient target variable modeling on very long sequences. We evaluate SHaRe-SSM on multiple such real-world datasets, ranging from 400 to almost 18k sequence lengths for classification. On average, SHaRe-SSM outperforms all first-order SSMs while being extremely energy efficient. Furthermore, we observe that SHaRe-SSM also outperforms first-order SSMs in an energy-efficient regression task on very long sequences of length 50k. In the future, we intend to work on Intel Loihi2 ShresthaLoihi2 implementation of SHaRe-SSM for edge AI applications.
@@ A.1 Background Theory
To study the theoretical properties of a second-order ODE in SHaRe-SSM, we take inspiration from [21]. Our model, just like their model, can be formulated as energy-conserving and with dissipative attributes.
A.1.1 Implicit Discretization (IM):
We consider the implicit (backward Euler) discretization of a second-order system involving a position-like state $u_{n}$ and a velocity-like state $v_{n}$, similar to [21]. The implicit scheme is known to introduce additional dissipative terms, which contribute to the stability of the dynamics, particularly in the presence of stiffness.
The discretized updates are given by:

### v2

@@ 3 Emperical Results
3.1 Human Activity Recognition
Our SNN model is capable of learning multiple states without any sequential recurrence, as it does not have a reset mechanism. Such a design enables parallelisation and reduced energy consumption. Hence, it is well-suited for wearable devices. Although our model is designed for long-range sequence datasets, we also evaluated its performance on Human Activity Recognition(HAR) datasets: UCI-HAR [anguita2013public] and SHAR datasets [micucci2017unimib]. For 30 subjects, UCI-HAR comprises 10.3k instances of 6 activities (walking, walking upstairs/downstairs, sitting, standing, lying) captured with a 3-axis accelerometer and gyroscope (50 Hz) on a Samsung Galaxy SII, while UniMB SHAR includes 11.7k instances of 17 activities (9 daily, 6 fall) recorded with a 3-axis accelerometer ($\leq 50$ Hz) on a Samsung Galaxy Nexus I9250. SHaRe-SSM achieves an accuracy of 99% and outperforms the best-performing model  [li2022wearablebased] for UCI-HAR by 0.2%. For SHAR, our model achieves 92.7%, which is better than non-spiking models in terms of energy and performance, and is competitive with spiking models, as it outperforms SpikeDCL by 0.6% and falls behind SpikeDCL by 1.2% (Refer to Table 1).
Model SNN UCIHAR SHAR
[…]
Table 1: Performance comparison (Accuracy %) between different networks on Human Activity Recognition datasets: UCI-HAR[anguita2013public], and SHAR [micucci2017unimib]
3.2 Very-Long Range Interactions
Wearable devices operate over long durations, demanding robust prediction on extended sequences. Spiking SSMs, with their high throughput and energy efficiency, are well-suited for this setting. To evaluate performance on very long temporal dependencies, we study the longest benchmark datasets: the Eigenworms dataset (17,984 sequences) for classification and the PPG-DaLiA dataset (49,920 sequences) for regression. Following the hyperparameter protocol of rusch2025oscillatory and employing Bayesian search [akiba2019optuna], we ensure an optimised training. Eigenworms, a subset of the UEA archive, represent C. elegans motion through six eigenworm features to distinguish wild-type behaviour from that of other mutants. PPG-DaLiA involves heart rate prediction from wrist-worn sensor data collected from 15 individuals over 150 minutes at 128 Hz.
For the Long-range classification task on EigenWorms, we observe that our model performs comparably to LinOSS with just $2.2\%$ drop in performance for IM discretisation, and significantly outperforms the IMEX counterpart by $10\%$. We observe an additional 82.6% energy improvement for the IMEX discretisation (Fig. 1). Our model outperforms all previous models, especially first-order SSMs such as LRU and S6 by $5\%$ and $7.8\%$ respectively.
We present the first regression results of a spiking SSM on extremely long sequences (up to 50k). To address the limited output range of spiking neurons, we introduce a kernel-based spiking regressor with a learnable temporal kernel. As shown in Table 2, SHaRe-SSM models consistently outperform all first-order SSMs. SHaRe-SSM-IMEX surpassing Mamba by $0.016$ MSE, demonstrating both the energy-efficient nature of IMEX discretisation and the strong representation power of resonating neurons. Despite their efficiency, our models are only at most $0.027$ MSE below second-order SSMs.
We observed from rusch2025oscillatory that IM discretisation outperforms IMEX for all datasets for classification and vice versa for regression. This trend is observed similarly for our model. Additionally, hu2024state demonstrated that linear SSM-based models outperform RNN-based models for chaotic systems. pourcel2025learning observed similar trends for the EigenWorms dataset, where the non-linear dynamics from unicornn suffer a tragic drop in performance; however, for the PPG dataset, it outperforms its linear counterpart, and only lags behind LinOSS-based models.
Method Integrator SNN EW (ACC $\uparrow$) PPG (MSE $\downarrow$)
NRDE[nrde] No 83.9 $\pm$ 7.3 9.9 $\pm$ 1.0
NCDE[ncde] No 75.0 $\pm$ 3.9 13.5 $\pm$ 0.7
Log-NCDE[walker2024log] No 85.6 $\pm$ 5.1 9.6 $\pm$ 0.6
LRU[lru] No 87.8 $\pm$ 2.8 12.2 $\pm$ 0.5
S5[s5] No 81.1 $\pm$ 3.7 12.6 $\pm$ 1.3
S6[mamba] No 85.0 $\pm$ 16.1 12.9 $\pm$ 2.1
Mamba[mamba] No 70.9 $\pm$ 15.8 10.7 $\pm$ 2.2
LinOSS[rusch2025oscillatory] IM No 95.0 $\pm$ 4.4 7.5 $\pm$ 0.5
Ours IM Yes 92.8 $\pm$ 3.3 $11.8\pm 0.9$
[…]
LinOSS[rusch2025oscillatory] IMEX No 80.0 $\pm$ 4.4 6.4 $\pm$ 0.2
Ours IMEX Yes 90.0 $\pm$ 5.7 9.1 $\pm$ 0.2
Table 2: Mean and Standard Deviation reported for longest sequences: Accuracy for EigenWorms, and Mean-Squared Error (MSE $\times 10^{-2}$) for PPG-DaLiA dataset across five training runs for the best model (Top three models are highlighted in red, blue, violet respectively).
@@ 3.2.1 Energy computation
Figure 1: Variation of energy computation of a LinOSS Block [rusch2025oscillatory] to the SHaRe-SSM Block for EigenWorms with respect to State (P), Hidden size (H).
@@
In this section, we assess and compare the energy computation for LinOSS and SHaRe-SSM for similar hidden-size(H), state-size(P), and num-blocks(N) on the EigenWorms dataset. From Figure 1, we observe that for different state-to-hidden ratios, we get different energy numbers. For the EigenWorms dataset, the average spike rate per time step per block is $0.42$. We observe our best performing IMEX model to be $\bf{\sim 52.1\times}$ energy efficient than an equivalent LinOSS model. Furthermore, since our model also outperforms LinOSS by 10%, it can captures state dynamics better than LinOSS.
@@ 4 Discussion
We propose SHaRe-SSM, an energy-efficient second-order spiking SSM built with harmonic resonate-and-fire neurons, along with a learnable encoder, decoder, and parallel scan method. Unlike prior similar SSMs, SHaRe-SSM is fully spike-based without GeLU/GLU and is tailored for very long-sequence modelling. The model performs well on HAR datasets while consuming significantly less energy, and also achieves superior classification and regression performance on 18k EigenWorms and 50k-length sequences PPG datasets, respectively. Hence, it is ideally suited for use in wearable devices in healthcare. Future work will focus on the deployment of SHaRe-SSM for edge AI using Intel Loihi2 [ShresthaLoihi2].
@@ A.1 Background Theory
To study the theoretical properties of a second-order ODE in SHaRe-SSM, we take inspiration from rusch2025oscillatory. Our model, like theirs, can be formulated as energy-conserving and possessing dissipative attributes.
A.1.1 Implicit discretisation (IM):
We consider the implicit (backward Euler) discretisation of a second-order system involving a position-like state $u_{n}$ and a velocity-like state $v_{n}$, similar to rusch2025oscillatory. The implicit scheme is known to introduce additional dissipative terms, which contribute to the stability of the dynamics, particularly in the presence of stiffness.
The discretised updates are given by: